# Dealing with Uncertainty in Contextual Anomaly Detection

### v1

@@ 4 Experiments
The goal of the empirical evaluation is to address the following three questions:
@@
  • RQ2: Does $i(x,y)$ correlate with sparsity in contextual feature space?
  • RQ3: Can NS aid in clinical practice for assessing the normalcy of the aortic diameter?
@@ 4.1 Data
Our experiments use four benchmark datasets and one real-world dataset.
Benchmark datasets. Table 1 summarizes the four UCI benchmark datasets, which include Abalone (Nash et al. 1994), containing physical measurements of abalones used to predict their age; Concrete Compressive Strength (Yeh 1998), describing the properties of concrete mixtures with the target being the compressive strength; Synchronous Machine (UCI 2021), representing the behavior of synchronous motors in electrical systems; and QSAR Fish Toxicity (Ballabio et al. 2015), which assesses the acute toxicity of chemicals on fish. Because these datasets do not contain contextual anomalies, we follow the experimental protocol introduced in (Li and Van Leeuwen 2023). Given a dataset $\{(x_{i},y_{i}),i=1,\dots,N\}$, where each context consists of $P$ variables, $x_{i}=\{x_{i1},\dots,x_{iP}\}$, the anomaly injection process modifies the dataset as follows:
@@
    For each selected instance $(x_{i},y_{i})$, perturb the behavioral feature $y_{i}$ to be $\hat{y}_{i}=y_{i}+\epsilon_{i}$, where $\epsilon_{i}$ is sampled uniformly from the range $[-0.5,-0.1]\cup[0.1,0.5]$. The contextual features $x_{i}$ remain unchanged which ensures that the anomaly arises purely due to the behavioral feature $\hat{y}_{i}$ deviating from its expected value given its context $x_{i}$.
@@
This scheme integrates the injected anomalies into the dataset while preserving the structure of the contextual features. Moreover, it addresses limitation of (Song et al. 2007)’s approach where some injected anomalies remained contextually normal. Table 1 shows the percent of perturbed examples in each dataset.
@@
Table 1: Summary of UCI benchmark datasets with injected anomalies.
Real-world dataset. We use a medical dataset on aorta normalcy (Frasconi et al. 2021). Aortic dilation is a risky condition with potentially deadly outcomes such as regurgitation or dissection. The dataset consists of 1110 normal subjects, 115 patients diagnosed with Marfan syndrome, and 351 patients with bicuspid aortic valve. The contextual variables are age, gender, weight, and height. We consider, separately, two behavioral variables, i.e. aorta diameters measured at two different locations: sinus of valsalva (SoV) and ascending aorta (AA).
@@ 4.2 Experimental Setup
We compare our approach to three types of state-of-the-art anomaly detections. First, we consider the contextual anomaly detection approaches QCAD and ROCOD as implemented in QCAD repository [^2] https://github.com/ZhongLIFR/QCAD. Second, we consider the classic Z-score. Finally, we use the standard non-contextual anomaly detection algorithms Isolation Forest (IForest) (Liu, Ting, and Zhou 2008), Local Outlier Factor (LOF) (Breunig et al. 2000), and Histogram-based Outlier Score (HBOS) (Goldstein and Dengel 2012) from the PyOD library (Zhao, Nasrullah, and Li 2019) with their default configurations.
The proposed NS relies on the heteroscedastic Gaussian Process Regressor (HGPR) model, implemented using the GPFlow library (Matthews et al. 2017). Key aspects of the HGPR model configuration include the use of a Rational Quadratic kernel (Rasmussen and Williams 2006), which is well-suited for capturing both smooth and non-smooth variations in the data. The heteroscedastic prior is modeled with a constant mean 0 and very low constant variance of $10^{-5}$. Inducing variables are initialized to the first 5% of the training data and updated during optimization, with the number of inducing points set proportionally to the dataset size to ensure scalability. The model is trained using a hybrid optimization approach that combines Natural Gradient Descent with $\gamma=0.02$ and the Adam optimizer with a learning rate of 0.01. We trained models for a maximum of 40,000 epochs.
To evaluate anomaly detection performance, we use a 5-fold cross-validation strategy. This ensures that every data point is evaluated as part of the test set once. We adopt widely accepted metrics in the anomaly detection literature (Liang and Parthasarathy 2016; Aggarwal 2017; Kuo, Li, and Kifer 2018; Li and Van Leeuwen 2023) like area under the ROC curve (ROC AUC), area under the precision recall curve (PR AUC), and Precision@n (which assesses the precision of the top-$n$ ranked points based on anomaly scores, where $n$ corresponds to the number of injected anomalies).
@@ 4.3 Results
RQ1: How does NS perform in detecting contextual anomalies compared to other methods? Table 2 reports the average over five different runs (repeating injection with different seeds) and standard deviation for each of the three evaluation metrics (ROC AUC, PR AUC, and Precision@n) across the benchmark datasets. The NS anomaly score is computed as the absolute value of the expected NS, as described in Section 3, ensuring that all deviations from normalcy are treated symmetrically, irrespective of direction.
The results demonstrate that NS achieves the best performance on three out of the four datasets, namely Abalone, SynMachine, and Toxicity, across most evaluation metrics. On the Abalone dataset, NS achieves the highest PR AUC (0.65 ± 0.04), highlighting its strong balance between precision and recall, which is particularly important for datasets where false positives have a significant impact. For the SynMachine dataset, NS shows near-perfect performance, with both ROC AUC and PR AUC reaching 1.00 ± 0.00, significantly outperforming other methods. Similarly, on the Toxicity dataset, NS achieves the highest PR AUC (0.67 ± 0.04) and a competitive ROC AUC (0.92 ± 0.02), showcasing its robustness in handling real-world challenges.
The Concrete dataset presents the only exception, where QCAD slightly outperforms NS in terms of ROC AUC and PR AUC.
@@
Non-contextual methods, including Isolation Forest, LOF, and HBOS, consistently underperform across all datasets and metrics. This result underscores the critical role of explicitly incorporating contextual information into anomaly detection models.
@@
Overall, the results highlight the robustness and versatility of NS, which consistently performs well across diverse domains and balances performance across multiple evaluation metrics.
Dataset Metric NS Z-score QCAD ROCOD IForest LOF HBOS
Abalone ROC AUC 0.96 ± 0.01 0.95 ± 0.01 0.90 ± 0.01 0.93 ± 0.01 0.77 ± 0.01 0.92 ± 0.00 0.67 ± 0.01
PR AUC 0.65 ± 0.04 0.57 ± 0.06 0.28 ± 0.04 0.40 ± 0.05 0.05 ± 0.01 0.39 ± 0.03 0.04 ± 0.00
P@n 0.61 ± 0.03 0.58 ± 0.04 0.39 ± 0.05 0.42 ± 0.06 0.03 ± 0.01 0.45 ± 0.04 0.03 ± 0.01
Concrete ROC AUC 0.89 ± 0.02 0.86 ± 0.03 0.93 ± 0.02 0.79 ± 0.02 0.62 ± 0.02 0.49 ± 0.06 0.74 ± 0.05
PR AUC 0.60 ± 0.01 0.55 ± 0.04 0.64 ± 0.04 0.34 ± 0.05 0.08 ± 0.02 0.06 ± 0.01 0.27 ± 0.05
P@n 0.57 ± 0.05 0.52 ± 0.05 0.62 ± 0.03 0.35 ± 0.05 0.08 ± 0.02 0.07 ± 0.04 0.32 ± 0.07
SynMachine ROC AUC 1.00 ± 0.00 1.00 ± 0.00 0.98 ± 0.01 0.90 ± 0.04 0.83 ± 0.03 0.92 ± 0.02 0.67 ± 0.03
PR AUC 1.00 ± 0.00 1.00 ± 0.00 0.96 ± 0.02 0.79 ± 0.08 0.33 ± 0.06 0.78 ± 0.04 0.26 ± 0.04
P@n 0.99 ± 0.02 1.00 ± 0.00 0.94 ± 0.03 0.73 ± 0.10 0.38 ± 0.06 0.71 ± 0.05 0.28 ± 0.05
Toxicity ROC AUC 0.92 ± 0.02 0.91 ± 0.02 0.86 ± 0.01 0.84 ± 0.03 0.71 ± 0.05 0.61 ± 0.05 0.69 ± 0.04
PR AUC 0.67 ± 0.04 0.57 ± 0.05 0.45 ± 0.07 0.46 ± 0.07 0.10 ± 0.02 0.08 ± 0.02 0.10 ± 0.02
P@n 0.62 ± 0.04 0.57 ± 0.06 0.46 ± 0.06 0.46 ± 0.07 0.09 ± 0.02 0.07 ± 0.03 0.10 ± 0.04
Table 2: Average ($\pm$ std) of the three considered metrics aggregated on a per-dataset basis. Our method obtains consistently the highest results for all metrics on Abalone and Toxicity, is comparable to Z-score on SynMachine, and ranks as second-best baseline on Concrete.
Model Sensitivity: Kernel and Inducing Points
@@
Table 4 illustrates the effect of increasing the number of inducing points from 5% to 20% of the training data. We observe that even with a small number of inducing points (5%), the model maintains reliable estimates of both mean and uncertainty, with only marginal benefits from further increases. This highlights a favorable trade-off between computational efficiency and predictive robustness, as the 5% configuration already achieves strong and consistent performance across datasets. However, training time increases significantly with more inducing points, scaling from approximately 10 minutes (5%) to around 70 minutes (20%), which may become a limiting factor in large-scale applications.
@@
Table 3: Impact of kernel choice on NS statistics.
@@
Table 4: Effect of inducing points ratio on NS and train time.
RQ2: Does $i(x,y)$ correlate with sparsity in contextual feature space?
We expect EU to concentrate in underrepresented (sparse) regions in the contextual space since the model has fewer training examples to rely on. In this experiment, we aim to verify that $i(x,y)$ does indeed align with sparse contextual regions. Specifically, we evaluate the ordinal association between $i(x,y)$ with the anomaly scores $s(x)$, computed on contextual variables only. Here $s(x)$ is obtained by applying three classic anomaly detectors: IForest, LOF and HBOS. The association is quantified by the Weighted Kendall Tau correlation coefficient (Shieh 1998). This rank-based metric is particularly suitable for comparing continuous variables, as it accounts for the magnitude of rank differences by assigning weights to the comparisons. In this version of the Kendall Tau coefficient, greater importance is given to elements higher in the rank. The coefficient is defined as:
@@
Table 5: Weighted Kendall Tau correlation between $i(x,y)$ and contextual anomaly scores $s(x)$.
To further explore the relationship, we visualize the distributions $i(x,y)$ for anomalous and inlier data points in the contextual space. Figure 2 shows kernel density estimation plots of $i(x,y)$ for the four datasets. In each case, the red curve represents the distribution for anomalous points (IForest), while the blue curve represents the distribution for inliers. It can be seen that anomalous points tend to have higher $i(x,y)$, reflecting higher uncertainty or sparsity in the contextual feature space.
(a) Abalone
(b) Concrete
(c) SynMachine
(d) Toxicity
Figure 2: KDE plots of $i(x,y)$ for anomalous and inlier points across all datasets. The red curves represent the distribution for anomalous points, while the blue curves represent the distribution for inliers.
@@ 4.4 Results on the Real-world Medical Data
RQ3: Can NS aid in clinical practice for assessing the normalcy of the aortic diameter?
For this experiment we train on normal subjects and test on adult patients with a diagnosis of Marfan or bicuspid aortic valve, two conditions typically associated with aorta dilations. We use the 40mm dilation threshold defined by current medical guidelines (Isselbacher et al. 2022) to assign patients to the positive class. Figure 3 shows the ROC for detecting anomalous SoV and AA aortic diameters, respectively. In both cases, our proposed approach results in a superior ROC curve and higher ROC AUC values (reported in the figure’s legend) than its contextual competitors. To verify the significance of the difference between the ROC curves, the Delong test (DeLong, DeLong, and Clarke-Pearson 1988) was applied, revealing statistically significant differences with QCAD and ROCOD ($p<0.001$) for both SoV and AA.
@@
One of the key advantages of NS is its ability to quantify EU via $i(x,y)$. Heatmaps in Figure 4 visualize $i(x,y)$ (males and females separately) at a fixed diameter of 32 mm (SoV and AA). To reduce the number of variables and enable 2D visualization, we use the body surface area (BSA, proportional to $\sqrt{W\times H}$) in the plots. Dark red regions indicate subpopulations with greater AU and correspond to regions of sparsity in the contextual space. This information can help clinicians identify and flag patients for whom the model’s predictions are less reliable, thus prioritizing additional diagnostic tests or closer monitoring. Such insights are especially valuable in clinical decision-making, where accurate and interpretable risk assessments are critical.
@@
Figure 4: Heatmaps of $i((\textrm{Age},\textrm{BSA}),y)$ at a fixed diameter $y=32$mm for SoV and AA diameters.
To further illustrate the utility of NS, we analyzed the dataset of 351 patients diagnosed with bicuspid aortic valve and deemed at risk of thoracic aortic aneurism. We first analyzed the different normalcy assessment of the traditional Z-score and NS. In Figure 5 we report results for each patient, after excluding 301 cases where both methods confirm the same result. Denoting by $(L,H)$ the HDI interval from NS, excluded cases are those where $Z>2$ and $L>2$ (both method confirm anomaly), and, symmetrically, cases where $Z<2$ and $H<2$ (both method confirm normality). In the figure, bars describe HDI intervals and are colored in red or in orange when the two assessments disagree. When the critical value 2 falls within the HDI, the normalcy assessment is unreliable and should be discarded (bars colored in blue).
Figure 5: Z-score vs. NS for normalcy of SoV diameters.
Figure 6 highlights six representative cases where the linear Z-score predicts an anomaly score below the clinical threshold $Z=2$ for at least one of the two aortic diameters (SoV or AA). However, the NS interval reveals significant diagnostic uncertainty, indicating that these patients may have a non-negligible probability of exceeding $Z=2$.
@@
(c) Patient 255
(d) Patient 265
(e) Patient 293
(f) Patient 297
Figure 6: Examples of bicuspid patients where the NS interval highlights cases with high diagnostic uncertainty. Each subplot represents a patient, with NS intervals shown as error bars.

### v2

@@ 4 Experiments
The goal of the empirical evaluation is to address the following four questions:
@@
  • RQ2: How sensitive is the framework to the choice of kernel and number of inducing points?
  • RQ3: Does $i(x,y)$ correlate with sparsity in contextual feature space?
  • RQ4: Can NS aid in clinical practice for assessing the normalcy of the aortic diameter?
@@ 4.1 Data
Benchmark datasets. Table 1 summarizes the five UCI benchmark datasets, which include Abalone (Nash et al., 1994), containing physical measurements of abalones used to predict their age; Concrete Compressive Strength (Yeh, 1998), describing the properties of concrete mixtures with the target being the compressive strength; Synchronous Machine (UCI, 2021), representing the behavior of synchronous motors in electrical systems; QSAR Fish Toxicity (Ballabio et al., 2015), which assesses the acute toxicity of chemicals on fish; and Yacht Hydrodynamics (Gerritsma et al., 1981), which involves predicting the hydrodynamic resistance of sailing yachts based on hull geometry and operating conditions.
Because these datasets do not contain contextual anomalies, we follow the experimental protocol introduced in Li and Van Leeuwen (2023) which injects anomalies into the entire dataset before performing any train/test split. Given a dataset $\{(x_{i},y_{i}),i=1,\dots,N\}$, where each context consists of $P$ variables, $x_{i}=\{x_{i1},\dots,x_{iP}\}$, the anomaly injection process modifies the dataset as follows:
@@
    For each selected instance $(x_{i},y_{i})$, perturb the behavioral feature $y_{i}$ to be $\hat{y}_{i}=y_{i}+\epsilon_{i}$, where $\epsilon_{i}$ is sampled uniformly from the range $0.5,-0.1]\cup[0.1,0.5]$. The contextual features $x_{i}$ remain unchanged which ensures that the anomaly arises purely due to the behavioral feature $\hat{y}_{i}$ deviating from its expected value given its context $x_{i}$.
@@
This scheme integrates the injected anomalies into the dataset while preserving the structure of the contextual features. Moreover, it addresses limitation of Song et al. (2007)’s approach where some injected anomalies remained contextually normal. Table 1 shows the percent of perturbed examples in each dataset.
Table 1: Summary of UCI benchmark datasets with injected anomalies.
@@
Yacht 308 9.7% Resistance
Real-world dataset. We use a medical dataset on aorta normalcy (Frasconi et al., 2021). Aortic dilation is a risky condition with potentially deadly outcomes such as regurgitation or dissection. The dataset consists of 1110 normal subjects, and 351 patients with bicuspid aortic valve. The contextual variables are age, gender, weight, and height. We consider, separately, two behavioral variables, i.e., aorta diameters measured at two different locations: sinus of valsalva (SoV) and ascending aorta (AA).
@@ 4.2 Experimental Setup
We compare our approach to three types of anomaly detection methods. First, we consider the contextual approaches QCAD and ROCOD, using the implementation from the QCAD repository [^2] https://github.com/ZhongLIFR/QCAD with the hyperparameter settings as in the original paper (number of trees in the quantile regression forest, maximal features, minimum node size, and number of conditional quantiles). Second, we include regression-based contextual baselines. For the Z-score in Equation 1, we fit a linear regression model on the training set and use a single global residual scale $S$ estimated from the training residuals. We further include the linear variance-estimation approach of Altman (1993) and, to isolate the contribution of heteroscedasticity, a homoscedastic variant of our method, denoted $\mathrm{NS}_{\mathrm{hom}}$, which uses a single GP. Finally, we use the standard non-contextual anomaly detection algorithms Isolation Forest (IForest) (Liu et al., 2008), Local Outlier Factor (LOF) (Breunig et al., 2000), and Histogram-based Outlier Score (HBOS) (Goldstein and Dengel, 2012) from the PyOD library (Zhao et al., 2019) with their default configurations.
The proposed NS relies on the heteroscedastic Gaussian Process Regressor (HGPR) model, implemented using the GPFlow library (Matthews et al., 2017). Key aspects of the HGPR model configuration include the use of a Rational Quadratic kernel (Rasmussen and Williams, 2006), which is well-suited for capturing both smooth and non-smooth variations in the data. The heteroscedastic prior is modeled with a constant mean 0 and very low constant variance of $10^{-5}$. Inducing variables are initialized to the first 5% of the training data and updated during optimization, with the number of inducing points set proportionally to the dataset size to ensure scalability. The model is trained using a hybrid optimization approach that combines Natural Gradient Descent with $\gamma=0.02$ and the Adam optimizer with a learning rate of 0.01. We trained models for a maximum of 40,000 epochs.
To evaluate anomaly detection performance, we use a 5-fold cross-validation strategy. This ensures that every data point is evaluated as part of the test set once. We adopt widely accepted metrics in the anomaly detection literature (Liang and Parthasarathy, 2016; Aggarwal, 2017; Kuo et al., 2018; Li and Van Leeuwen, 2023) like area under the ROC curve (ROC AUC) and area under the precision recall curve (PR AUC).
@@ 4.3 Results
RQ1: How does NS perform in detecting contextual anomalies compared to other methods? Table 2 reports the average over five independent runs (each repeating the injection procedure with a different seed) together with the standard deviation for both metrics across the benchmark datasets. The NS anomaly score is computed as the absolute value of the expected NS, as detailed in Section 3.2, so that deviations in either direction are treated symmetrically.
NS secures the top ROC AUC and PR AUC on Abalone, SynMachine, Toxicity, and Yacht, with Concrete remaining the lone dataset where QCAD holds a narrow edge. NS attains perfect scores on SynMachine; its strong $0.97/0.88$ ROC AUC/PR AUC on Yacht, together with competitive PR AUC values of $0.65$ on Abalone and $0.67$ on Toxicity, underscores the method’s robustness across heterogeneous datasets.
Both $\mathrm{NS}_{\mathrm{hom}}$ and the Altman method deliver strong results on several datasets, while NS remains the most consistently strong approach overall. Nonetheless, it is striking that an off‑the‑shelf Z‑score baseline, overlooked in prior work, already proves remarkably competitive on this benchmark. To the best of our knowledge, this is the first work to include such a Z‑score baseline in a contextual anomaly detection benchmark, and its strong results show it should be treated as a meaningful reference point.
@@
Non-contextual methods, including Isolation Forest, LOF, and HBOS, consistently underperform across all datasets and metrics. This result underscores the critical role of explicitly incorporating contextual information into anomaly detection models. Importantly, the gap is driven by the problem formulation. These methods model $p(x,y)$ rather than the conditional $p(y\mid x)$, and the effect is not due to the particular choice of detector. We therefore use standard non-contextual baselines commonly reported in prior CAD studies.
@@
Overall, the results highlight the robustness and versatility of NS, which consistently performs well across diverse domains and balances performance across diverse evaluation metrics.
Table 2: Average ($\pm$ std) ROC AUC and PR AUC computed across five independent anomaly injections (each evaluated with 5-fold cross-validation).
Abalone Concrete SynMachine Toxicity Yacht
Method ROC AUC PR AUC ROC AUC PR AUC ROC AUC PR AUC ROC AUC PR AUC ROC AUC PR AUC
NS 0.96 ± 0.01 0.65 ± 0.04 0.89 ± 0.02 0.60 ± 0.01 1.00 ± 0.00 1.00 ± 0.00 0.92 ± 0.02 0.67 ± 0.04 0.97 ± 0.02 0.88 ± 0.06
$\mathrm{NS}_{hom}$ 0.95 ± 0.03 0.64 ± 0.06 0.86 ± 0.02 0.52 ± 0.03 1.00 ± 0.00 1.00 ± 0.00 0.92 ± 0.02 0.63 ± 0.05 0.95 ± 0.01 0.57 ± 0.06
Z-score 0.95 ± 0.01 0.57 ± 0.06 0.86 ± 0.03 0.55 ± 0.04 1.00 ± 0.00 1.00 ± 0.00 0.91 ± 0.02 0.57 ± 0.05 0.82 ± 0.04 0.53 ± 0.09
Altman 0.95 ± 0.01 0.48 ± 0.05 0.87 ± 0.03 0.58 ± 0.06 0.99 ± 0.01 0.96 ± 0.03 0.91 ± 0.02 0.61 ± 0.03 0.81 ± 0.05 0.59 ± 0.11
QCAD 0.90 ± 0.01 0.28 ± 0.04 0.93 ± 0.02 0.64 ± 0.04 0.98 ± 0.01 0.96 ± 0.02 0.86 ± 0.01 0.45 ± 0.07 0.96 ± 0.03 0.85 ± 0.11
ROCOD 0.93 ± 0.01 0.40 ± 0.05 0.79 ± 0.02 0.34 ± 0.05 0.90 ± 0.04 0.79 ± 0.08 0.84 ± 0.03 0.46 ± 0.07 0.78 ± 0.03 0.31 ± 0.05
IForest 0.77 ± 0.01 0.05 ± 0.01 0.62 ± 0.02 0.08 ± 0.02 0.83 ± 0.03 0.33 ± 0.06 0.71 ± 0.05 0.10 ± 0.02 0.79 ± 0.06 0.32 ± 0.10
LOF 0.92 ± 0.00 0.39 ± 0.03 0.49 ± 0.06 0.06 ± 0.01 0.92 ± 0.02 0.78 ± 0.04 0.61 ± 0.05 0.08 ± 0.02 0.71 ± 0.04 0.20 ± 0.04
HBOS 0.67 ± 0.01 0.04 ± 0.00 0.74 ± 0.05 0.27 ± 0.05 0.67 ± 0.03 0.26 ± 0.04 0.69 ± 0.04 0.10 ± 0.02 0.76 ± 0.05 0.34 ± 0.08
Assessing the reliability of HDI We evaluate using NS in a learning with abstention setting, where a model only offers a prediction on the test examples for which it is most confident. By abstaining when the model is uncertain, its performance on those examples for which it does offer a prediction should improve. We consider two different approaches for abstention when using NS. First, we use the HDI length $i(x,y)$: we sort all test points by their HDI length and abstain on the 5% with the widest HDI. Second, we fit an Isolation Forest on the contextual variables and abstain on the test points that have the top 5% most “outlying” contexts. Table 3 reports the ROC AUC and PR AUC computed over the 95% most confident examples identified by each approach. Across datasets the HDI-based abstention strategy yields superior or equal performance compared to the contextual baseline, supporting the use of $i(x,y)$ as a reliable indicator of epistemic uncertainty.
Table 3: ROC and PR AUC when making predictions on the 95% most certain test instances as determined by i(x,y) and IForest on contextual space when using the NS.
Dataset Abstain using $i(x,y)$ Abstain using IForest
ROC AUC PR AUC ROC AUC PR AUC
Abalone 0.97 0.71 0.96 0.66
Concrete 0.92 0.65 0.86 0.55
SynMachine 1.00 1.00 1.00 1.00
Toxicity 0.95 0.74 0.92 0.70
Yacht 1.00 0.95 0.97 0.87
RQ2: How sensitive is the framework to the choice of kernel and number of inducing points?
@@
Table 5 illustrates the effect of increasing the number of inducing points from 5% to 20% of the training data. We observe that even with a small number of inducing points (5%), the model maintains reliable estimates of both mean and uncertainty, with only marginal benefits from further increases. This highlights a favorable trade-off between computational efficiency and predictive robustness, as the 5% configuration already achieves strong and consistent performance across datasets. However, training time increases significantly with more inducing points, scaling from approximately 10 minutes (5%) to around 70 minutes (20%) on an NVIDIA A6000 GPU.
Table 4: Impact of kernel choice on NS statistics.
@@
Table 5: Effect of inducing points ratio on NS and train time.
@@
RQ3: Does $i(x,y)$ correlate with sparsity in contextual feature space?
We expect EU to concentrate in underrepresented (sparse) regions in the contextual space since the model has fewer training examples to rely on. In this experiment, we aim to verify that $i(x,y)$ does indeed align with sparse contextual regions. Specifically, we evaluate the ordinal association between $i(x,y)$ with the anomaly scores $s(x)$, computed on contextual variables only. Here $s(x)$ is obtained by applying three classic anomaly detectors: IForest, LOF and HBOS. The association is quantified by the Weighted Kendall Tau correlation coefficient (Shieh, 1998). This rank-based metric is particularly suitable for comparing continuous variables, as it accounts for the magnitude of rank differences by assigning weights to the comparisons. In this version of the Kendall Tau coefficient, greater importance is given to elements higher in the rank. The coefficient is defined as:
@@
Table 6: Weighted Kendall Tau correlation between $i(x,y)$ and anomaly scores $s(x)$.
@@
Yacht 0.68 0.70 0.65
@@ 4.4 Results on the Real-world Medical Data
RQ4: Can NS aid in clinical practice for assessing the normalcy of the aortic diameter?
For this experiment we train on normal subjects and test on adult patients with a diagnosis of bicuspid aortic valve, a condition typically associated with aorta dilations. We use the 40mm dilation threshold defined by current medical guidelines (Isselbacher et al., 2022) to assign patients to the positive class. This threshold is used solely to determine the ground truth labels needed for the ROC analysis and qualitative interpretation. NS assesses normalcy as a continuous quantity conditioned on patient context and is not trained on or tuned to this 40 mm cut-off.
Figure 2 shows the ROC for detecting anomalous SoV and AA aortic diameters, respectively. In both cases, our proposed approach results in a superior ROC curve and higher ROC AUC values (reported in the figure’s legend) than its contextual competitors. To verify the significance of the difference between the ROC curves, the Delong test (DeLong et al., 1988) was applied, revealing statistically significant differences with QCAD and ROCOD ($p<0.001$) for both SoV and AA.
@@
One of the key advantages of NS is its ability to quantify EU via $i(x,y)$. Heatmaps in Figure 3 visualize $i(x,y)$ (males and females separately) at a fixed diameter of 32 mm (SoV and AA). To reduce the number of variables and enable 2D visualization, we use the body surface area (BSA, proportional to $\sqrt{W\times H}$) in the plots. Dark red regions indicate subpopulations with greater EU and correspond to regions of sparsity in the contextual space. This information can help clinicians identify and flag patients for whom the model’s predictions are less reliable, thus prioritizing additional diagnostic tests or closer monitoring. Such insights are especially valuable in clinical decision-making, where accurate and interpretable risk assessments are critical.
@@
Figure 3: Heatmaps of $i((\textrm{Age},\textrm{BSA}),y)$ at a fixed diameter y=32y=32mm for SoV and AA diameters.
Figure 4 highlights four representative cases where the linear Z-score predicts an anomaly score below the clinical threshold $Z=2$ for at least one of the two aortic diameters (SoV or AA). However, the NS interval reveals significant diagnostic uncertainty, indicating that these patients may have a non-negligible probability of exceeding $Z=2$.
@@
(c) Patient 293
(d) Patient 297
Figure 4: Bicuspid patients where NS interval highlights high diagnostic uncertainty. For each diameter, we show the Z-score (red), NS (blue), its HDI (error bar), and the corresponding probability $P(\mathrm{NS}>2)$.
In clinical practice, it may also be intuitive to report the posterior probability that NS exceeds a clinically meaningful threshold, rather than inspecting the full high-density interval. For the bicuspid cohort, we therefore compute the exceedance probability $P(\mathrm{NS}(x,y)>2)$ for each patient and report it alongside the per-diameter intervals in Figure 4. This summary can help flag borderline cases beyond point estimates: among patients with $\mathbb{E}[\mathrm{NS}]<2$, 82 (SoV) and 48 (AA) still have $P(\mathrm{NS}>2)>0$ (i.e., a non-zero probability of crossing the anomaly threshold).